Earlier reports Convert Day of Year

The Hubble Space Telescope

Unlocking the Secrets of the Universe

Daily Report #5812

Period Covered:
08:00 pm June 11, 2012 - 07:59 pm June 12, 2012
(DOY 164/0000z - 164/2359z)

Flight Operations Summary

Significant Spacecraft Anomalies: (The following are preliminary reports
of potential non-nominal performance that will be investigated.)


19271-0 - Clear COS Event Flag 3@164/1527z


FGS GSAcq		9		9
FGS REAcq		8		8
OBAD with Maneuver	6		6

 The COS FUV Special Recovery activities (proposal 12810 visits 26 - 27) successfully completed 
today without incident.  The FUV was ramped to -5293V for segment A (178 counts) and -5246V (175 
count) for B, these are the "normal" HV levels.  The QE grid was on.   No elevated counts or HV 
current transients were observed in the engineering telemetry during or after the ramp.  The STScI 
and MOSES2 Engineering Teams have examined the relevant telemetry and confirmed all voltages, 
currents and temperatures were as expected. The voltage and current monitors matched the signatures 
seen in SMOV.  The COS Science Team at STScI has examined the science data and DCE memory dumps and 
everything is nominal.  No HV current transients or unexpected elevated count rates were observed 
in the DCE memory dumps. The science data consisted of 2 sets of dark and wavecal exposures. The 
exposures were collected immediately after ramping to HVNom (178/175).  Comparisons with SMOV data 
taken indicates comparable behavior when the aging of the detector and temporal fluctuations are 
considered.  The Engineering and Science Teams have given a “Go” to resume normal FUV HV 
activities.  The COS event flag 3 was cleared @164/15:30 via OR 19271.  Many thanks to all who help 
investigate, test, plan and successfully execute the FUV HV Recovery.  Next FUV HV activity is 
currently schedule for 165/13:10 as part of proposal 12716 (FUV Detector Dark Monitor)  

Program Principal Investigator Program Title
Science Observations Scheduled
12457 Marc Postman, Space Telescope Science Institute Through a Lens, Darkly - New Constraints on the Fundamental Components of the Cosmos
12521 Xin Liu, University of California - Los Angeles The Frequency and Demographics of Dual Active Galactic Nuclei
12528 Philip Massey, Lowell Observatory Probing the Nature of LBVs in M31 and M33: Blasts from the Past
12549 Thomas Brown, Space Telescope Science Institute The Formation History of the Ultra-Faint Dwarf Galaxies
12668 Slawomir Piatek, New Jersey Institute of Technology Proper Motion Survey of Classical and SDSS Local Group Dwarf Galaxies
12801 Harold Weaver, The Johns Hopkins University Applied Physics Laboratory Hubble Deep Search for Debris and Satellites in the Pluto System in Support of NASA's New Horizons Mission
Calibration Observations Scheduled
12688 Tiffany Borders, Space Telescope Science Institute Cycle 17: UVIS Bowtie Monitor
12689 Tiffany Borders, Space Telescope Science Institute WFC3 UVIS CCD Daily Monitor
12742 Justin Ely, Space Telescope Science Institute CCD Dark Monitor Part 2
12744 Justin Ely, Space Telescope Science Institute CCD Bias and Read Noise Monitor-Part 2
12765 Justin Ely, Space Telescope Science Institute STIS CCD Hot Pixel Annealing
12766 Elena Mason, Space Telescope Science Institute - ESA STIS CCD Spectroscopic Flat-field Monitor
12784 Elena Sabbi, Space Telescope Science Institute - ESA Characterization of UVIS Traps Via Charge-Injected Biases
12803 Knox Long, Space Telescope Science Institute Guard Darks
Other Observations Scheduled
12810 Thomas Wheeler, Space Telescope Science Institute COS/FUV Special Recovery